A silver dollar is flipped twice. Calculate the proability of each of the following occurring:
a) a head on the first flip
b) a tail on the second flip given that the first toss was a head
c) two tails
d) a tail on the first and a head on the second
e) a tail on the first and a head on the second or a head on the first and a tail on the second
e) at least one head on the two flips

Found 2 solutions by stanbon, Fombitz:
Answer by stanbon(75887)

A silver dollar is flipped twice. Calculate the proability of each of the following occurring:
Note: The flip results are independent and P(H)=P(T)=1/2
--------------------------------------
a) a head on the first flip:: 1/2
------------------------------------------
b) a tail on the second flip given that the first toss was a head::1/2
----
c) two tails:::(1/2)(1/2) = 1/4
-------------------------------------------
d) a tail on the first and a head on the second: (1/2)(1/2)= 1/4
-------------------------------------------
e) a tail on the first and a head on the second or a head on the first and a tail on the second:::(1/2)(1/2) + (1/2)(1/2) = 2/4 = 1/2
---------------------------------------------------
e) at least one head on the two flips::: 1 - P(no heads) = 1-(1/4) = 3/4
==============================================================
Cheers,
Stan H.
===============

Answer by Fombitz(32388)

a) Look at the 2 possible outcomes.
H
T
P(H)=1/2
.
.
b) Look at the 2 possible outcomes.
H H
H T
P(T)=1/2
.
.
c)Look at the 4 possible outcomes.
H H
H T
T H
T T
P(T,T)=1/4
.
.
d)Look at the 4 possible outcomes.
H H
H T
T H
T T
P(T,H)=1/4
.
.
e) From d) P(T,H)=1/4
P(H,T)=1/4
P(T,H or H,T)=1/4+1/4=1/2
.
.
f)
P(at least one head)+P(all tails)=1
P(at least one head)=1-P(T,T)
P(at least one head)=1-1/4
P(at least one head)=3/4